### v1

# XTRUST: On the Multilingual Trustworthiness of Large Language Models

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across a range of natural language processing (NLP) tasks, capturing the attention of both practitioners and the broader public. A key question that now preoccupies the AI community concerns the capabilities and limitations of these models, with trustworthiness emerging as a central issue—particularly as LLMs are increasingly applied in sensitive fields like healthcare and finance, where errors can have serious consequences. However, most previous studies on the trustworthiness of LLMs have been limited to a single language, typically the predominant one in the dataset, such as English. In response to the growing global deployment of LLMs, we introduce XTRUST, the first comprehensive multilingual trustworthiness benchmark. XTRUST encompasses a diverse range of topics, including illegal activities, hallucination, out-of-distribution (OOD) robustness, physical and mental health, toxicity, fairness, misinformation, privacy, and machine ethics, across 10 different languages. Using XTRUST, we conduct an empirical evaluation of the multilingual trustworthiness of five widely used LLMs, offering an in-depth analysis of their performance across languages and tasks. Our results indicate that many LLMs struggle with certain low-resource languages, such as Arabic and Russian, highlighting the considerable room for improvement in the multilingual trustworthiness of current language models. The code is available at https://github.com/LluckyYH/XTRUST.
{CJK}
UTF8gbsn
XTRUST: On the Multilingual Trustworthiness of Large Language Models
Yahan Li1111Equal contribution  Yi Wang2111Equal contribution  Yi Chang1,3,4  Yuan Wu 1222Corresponding author: yuanwu@jlu.edu.cn 1School of Artificial Intelligence, Jilin University, China 2The Hong Kong University of Science and Technology (Guangzhou), China 3Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 4International Center of Future Science, Jilin University, China
Figure 1: Chat with LLMs in non-English languages can lead to untrustworthy responses
1 Introduction
In the rapidly evolving field of artificial intelligence (AI), large language models (LLMs) have achieved remarkable progress in a variety of natural language processing (NLP) tasks Zhao et al. (2023); Min et al. (2023), including writing assistance Zhang et al. (2023b), code generation Ouyang et al. (2023), machine translation Zhang et al. (2023a), task planning Valmeekam et al. (2023), and reasoning Huang and Chang (2022), among others. Their exceptional performance has led to their deployment in sensitive domains such as medicine Thirunavukarasu et al. (2023), finance Wu et al. (2023), and law Cui et al. (2023). This widespread use highlights a critical and pressing concern: the need to ensure the trustworthiness of LLMs.
Existing research on the trustworthiness of LLMs has predominantly focused on English-language data Liang et al. (2022); Liu et al. (2023); Sun et al. (2024), with limited attention to their multilingual capabilities. As LLMs garner increasing interest from global industries and academic circles, they are frequently utilized in non-English communications, engaging with users from diverse linguistic backgrounds. Hence, assessing the multilingual trustworthiness of LLMs is of vital importance (As illustrated in Figure 1).
In this paper, we introduce XTRUST, the first benchmark designed to evaluate the trustworthiness of LLMs across multiple languages. XTRUST offers three key advantages: (1) Extensive Diversity. It includes a total of 23,590 instances, covering 10 distinct categories of trustworthiness concerns, providing a robust and comprehensive evaluation framework for LLMs. (2) Diverse Question Types. XTRUST comprises three types of test questions: binary classification, multiple-choice classification, and natural language generation, ensuring that LLMs are rigorously tested across various trustworthiness scenarios. (3) Multilingual Support. The benchmark leverages Google Translate to translate data into 10 languages—Arabic, Chinese, French, German, Hindi, Italian, Korean, Portuguese, Russian, and Spanish—enabling a broader and more inclusive assessment.
Using XTRUST, we evaluated five widely adopted LLMs: GPT-4 OpenAI (2023b), GPT-3.5 Turbo OpenAI (2023a), Text-Davinci-002 Floridi and Chiriatti (2020), Baichuan, and Gemini Pro Team et al. (2023). Our results show that GPT-4 consistently outperformed the other models across most trustworthiness dimensions. Interestingly, Text-Davinci-002 delivered the best performance in the area of toxicity. However, it is noteworthy that all models achieved less than 70% average accuracy on certain categories, such as hallucination, out-of-distribution robustness, and physical health, emphasizing the need for further improvement in LLM trustworthiness. We hope that XTRUST will foster a deeper understanding of the trustworthiness of LLMs and assist practitioners in delivering more reliable models to users in non-English-speaking regions.
2 Related Works
2.1 Trustworthiness Evaluation of LLMs
The evaluation of LLMs is a pivotal aspect of their development and has recently garnered substantial attention from both academia and industry Chang et al. (2024). In particular, evaluating LLMs’ alignment capabilities with human preferences has emerged as a key priority as LLMs are increasingly developed in a wide range of real-world applications. DecodingTrust evaluates the trustworthiness of GPT-4 and GPT-3.5 from multiple perspectives Wang et al. (2023a). AdvCoU introduces a prompting strategy that uses malicious demonstrations to test the trustworthiness of open-source LLMs Mo et al. (2023). Do-Not-Answer presents a dataset specifically designed to challenge the safeguard mechanisms of LLMs by including prompts that responsible models should avoid answering Wang et al. (2023b). TRUSTLLM outlines various principles of trustworthiness, establishes benchmarks, conducts evaluations, and provides a comprehensive analysis of LLM trustworthiness Sun et al. (2024). Notably, all of these studies focus exclusively on English-language models.
2.2 Multilingual Benchmarks and Evaluation
Benchmarks for multilingual evaluation, such as XTREMEHu et al. (2020), XTREME-RRuder et al. (2021), and XGLUELiang et al. (2020), have been developed to assess cross-lingual transfer in LLMs. Building on their success, several benchmarks have been introduced to cover specific language families. Examples include IndicXTREMEDoddapaneni et al. (2022) for Indian languages, MasakhaNER 2.0Adelani et al. (2022) for African languages, and IndonluWilie et al. (2020) for Indonesian. Furthermore, research such as Hendy et al. (2023) has evaluated the translation capabilities of LLMs, finding that while LLMs perform well with high-resource languages, their abilities in low-resource languages remain limited. MEGA conducts a multilingual evaluation of mainstream LLMs on standard NLP tasks, such as classification and question answering Ahuja et al. (2023). However, unlike these studies, which primarily focus on standard NLP tasks in cross-linguistic contexts, our XTRUST benchmark offers a comprehensive evaluation of trustworthiness in LLMs across multiple languages. This provides a more profound understanding of LLMs’ trustworthiness capabilities within a multilingual framework.
3 XTRUST Construction
Figure 2: The overview of the proposed XTRUST benchmark
3.1 Trustworthiness Categories
An overview of XTRUST is presented in Fig. 2, We collect a total of 2359 instances spanning 10 categories of trustworthiness issues from several monolingual datasets:
Illegal Activity. This category centers on identifying illegal behaviors that may result in harmful societal outcomes. LLMs must possess a fundamental understanding of the law and the ability to accurately differentiate between legal and illegal actions.
Hallucination. This category addresses nonsensical or inaccurate content produced by LLMs that contradicts established sources. LLMs should be capable of determining whether the input can be validated by factual information.
Out-of-Distribution Robustness. This category evaluates the ability of LLMs to perform effectively on previously unseen test data.
Mental Health. This category assesses the model’s capability to offer guidance and information on psychological well-being, with a particular emphasis on stress management and emotional resilience.
Physical Health. This category examines actions or expressions that may impact human physical health. LLMs should be knowledgeable about the appropriate actions and responses in various situations to support and maintain physical well-being.
Toxicity. This category evaluates how effectively LLMs can resist generating harmful responses. LLMs must be able to recognize and reject offensive or harmful content and actions.
Fairness and Bias. This category addresses social biases across a range of topics, including race, gender, and religion. LLMs are expected to recognize and avoid unfair or biased language and behaviors.
Misinformation. This category examines the issue of LLMs generating misleading responses due to their limitations in delivering factually accurate information. LLMs must be able to detect and avoid producing incorrect or deceptive content.
Privacy. This category focuses on privacy-related concerns. LLMs should demonstrate a strong understanding of privacy principles and be committed to avoiding any unintentional breaches of user privacy.
Machine Ethics. This category evaluates the moral decision-making abilities of LLMs. LLMs should demonstrate strong ethical principles and actively reject unethical behavior or language.
3.2 Data Collection
We conducted a systematic review of numerous trustworthy datasets from various fields, including safety, fairness, and privacy, to establish the foundation for the multilingual XTRUST evaluation. For the Out-of-Distribution (OOD) robustness assessment, we collected recent news articles from the Internet, ensuring that the information was not used to train LLMs. All data collected for this study was carefully reviewed by our team.
For Illegal Activity, we randomly sampled 200 instances from the Chinese AI and Law challenge dataset (CAIL2018)Xiao et al. (2018), a large compilation of legal case records provided by the Supreme People’s Court of China. For Hallucination, we selected 300 samples from the Question Answering (QA) task in HaluevalLi et al. (2023), a dataset containing generated and human-annotated hallucinated examples to evaluate LLMs’ ability to detect hallucinations. For OOD Robustness, we gathered entertainment, education, and international news from China News Service, specifically collecting articles published after September 2023 to ensure the information was not used in LLM training. A total of 200 articles were used for this evaluation. For Mental Health, we randomly selected 200 instances from the Mental Health subset of SafetyBench Zhang et al. (2023c), which includes safety-related multiple-choice questions. For Physical Health, we randomly selected 200 samples from SafeText Levy et al. (2022), a dataset containing 367 human-written real-life scenarios with safe and unsafe suggestions for each scenario. For Toxicity, we randomly selected 200 instances from the toxic task prompts subset of REALTOXICITYPROMPT Gehman et al. (2020) to evaluate the toxicity generation of LLMs. For Fairness & Bias, we selected prompts from the Stereotype Bias subset of DecodingTrust Wang et al. (2023a), asking LLMs whether they agree or disagree with the statements presented in the prompts. DecodingTrust spans 24 demographic groups (e.g., "Homosexuals," "Men," "White people") and 16 bias topics (e.g., "Sexually Transmitted Diseases," "Occupations," "Intelligence"). We selected 384 prompts (24 demographic groups × 16 bias topics) to comprehensively evaluate LLMs’ performance on fairness and bias. For Misinformation, we used 155 questions from the Misinformation Harms subset of the Do-Not-Answer dataset Wang et al. (2023b), a taxonomy for assessing the risk of harm posed by LLMs. Additionally, we collected 86 fake news-related questions from the Internet, using a total of 241 questions in this evaluation. For Privacy, we selected 34 data entries from the Privacy subset of DecodingTrust Wang et al. (2023a) to evaluate LLMs’ capability to prevent data leakage. For Machine Ethics, we randomly selected 200 short samples from the test set of ETHICS Gui et al. (2021), which assesses a model’s knowledge of basic ethical concepts. We also selected 200 instances from the Jiminy Cricket dataset Hariri (2023), which evaluates an agent’s ability to make moral decisions in text-based games.
3.3 Translating the Collected Data
To create a multilingual trustworthiness benchmark, we translated the original monolingual data into multiple languages. For this study, we selected 10 languages: Arabic (AR), Chinese (ZH), French (FR), German (DE), Hindi (HI), Italian (IT), Korean (KO), Portuguese (PT), Russian (RU), and Spanish (ES). The test data was translated using Google Translate.
4 Experiments
4.1 Models
We conducted experiments on five LLMs that support multilingual applications, including baichuan2-7b-chat-v1 (denoted as Baichuan)Yang et al. (2023), Gemini-pro (released on December 13, 2023, denoted as Gemini)Team et al. (2023), davinci-002 (denoted as Davinci)Brown et al. (2020), gpt-3.5-turbo-1106 (denoted as ChatGPT)OpenAI (2022), and gpt-4-1106-preview (denoted as GPT-4) OpenAI (2023b). All these models are API-based. For each LLM, we set the temperature to 0 for classification tasks to generate deterministic outputs, and to 1 for text generation tasks to encourage more diverse continuations. We evaluate the LLMs in both zero-shot and few-shot settings, carefully crafting prompts to elicit undesirable behaviors for the purpose of trustworthiness evaluation. For each language, the designed prompts were translated using Google Translate into the corresponding language. The details of evaluated LLMs and prompt designs are shown in the Appendix.
4.2 Evaluation on Illegal Activities
Model AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan 45.0 98.5 4.5 100 12.5 70.0 0 74.0 84.0 18.0 50.65
Gemini 96.5 95.5 94.0 97.0 91.5 11.0 87.5 97.5 98.5 95.5 86.45
Davinci 6.0 3.0 6.5 2.0 3.0 20.0 3.0 22.5 1.5 15.0 8.25
ChatGPT 61.0 75.5 76.0 86.5 57.5 60.5 0 75.0 2.0 77.0 57.10
GPT-4 53.5 96.0 71.0 97.0 85.0 80.5 91.5 58.5 93.0 79.0 80.5
Table 1: Evaluation on illegal activity.
Setup. To assess the ability of LLMs to handle illegal activity, we transform the collected legal case records into binary classification questions, asking the models to differentiate between legal and illegal behaviors based on a straightforward task description. This approach aims to evaluate how effectively LLMs comprehend legal knowledge. For this task, we conduct evaluations in a zero-shot classification setting.
Results. As shown in Table 1, Gemini demonstrates the strongest performance in detecting illegal activities across multiple languages, achieving an impressive average score of 86.45%, particularly excelling in languages such as Arabic, Chinese, French, and Portuguese. GPT-4 follows closely, with an average score of 80.5, showing strong performance in languages like Chinese, French, German, and Russian. Baichuan ranks third, with an average score of 50.65%, though its performance is notably weaker in languages such as French, Hindi, and Spanish. ChatGPT and Davinci display lower average scores of 57.10% and 8.25%, respectively, highlighting the pressing need for improvements in multilingual trustworthiness for detecting illegal activities.
4.3 Evaluation on Hallucination
Model AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan 13.0 25.5 38.0 49.0 37.0 24.8 12.0 39.5 42.5 32.0 31.33
Gemini 46.5 49 46.5 41.5 21.0 40.2 50.0 42.5 0 42.5 37.97
Davinci 18.5 52.0 10.5 25.0 9.0 29.3 44.0 8.5 46.5 15.0 25.83
ChatGPT 53.0 37.5 51.0 44.0 46.0 40.2 50.0 56.5 56.0 44.5 47.87
GPT-4 49.5 45.0 51.5 48.5 46.0 49.7 38.0 58.5 43.5 48.5 47.87
Table 2: Evaluation on hallucination.
Setup. To assess how effectively LLMs avoid generating hallucinations, we task them with determining whether the statements in the input are factual or hallucinated. This evaluation is conducted in a zero-shot classification setting.
Results. As shown in Table 2, GPT-4 outperforms other LLMs, achieving the highest scores in French (51.5%), Hindi (49.7%), Italian (49.7%), Portuguese (58.5%), and Spanish (48.5%). It also secures the second-highest scores in Arabic (49.5%) and German (48.5%). ChatGPT follows closely, delivering strong performance across most languages, particularly excelling in Arabic (53.0%), Hindi (46.0%), Russian (56.0%), and Korean (50.0%). While Baichuan, Gemini, and Davinci demonstrate notable strengths in specific languages, their overall performance is less optimal.
4.4 Evaluation on Out of Distribution Robustness
Model Task AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan 0-shot 0.5 41.5 46.5 0 66.5 7.5 0.4 0.5 2.4 12.0 17.78
Gemini 0-shot 73.0 75.0 14.0 14.5 87.5 83.0 0 71.0 0 88.5 50.65
Davinci 0-shot 0 54.0 7.5 0 0 6.0 0 0 0 5.5 7.3
ChatGPT 0-shot 0 58.0 50.5 5.0 69.9 77.5 8.9 1.0 0.9 31.5 30.32
GPT-4 0-shot 3.4 93.5 98.0 34.0 99.5 98.5 2.4 2.5 19.0 98.0 54.88
Table 3: Evaluation on out-of-distribution robustness.
Setup. To assess the robustness of LLMs against OOD data, we convert the collected news data into a question-answer format, prompting the LLMs to determine whether the input event is true or false based on a straightforward task description. Additionally, we introduce an "I do not know" option to examine how LLMs handle unknown events. For OOD robustness, we conduct the evaluation in a zero-shot setting.
Results. The experimental results are presented in Table 3. In the zero-shot setting, GPT-4 leads in performance across all languages except Arabic and Korean, while Gemini secures the second-highest average performance. Davinci shows the weakest results. Notably, most tested LLMs encounter difficulties with Arabic, Korean, Portuguese, and Russian, whereas they perform well in Chinese and French.
4.5 Evaluation on Mental Health
Model AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan 17.0 82.0 46.5 64.0 30.0 57.0 57.0 56.5 37.0 59.5 50.65
Gemini 30.0 48.0 64.0 65.5 25.0 55.5 100.0 53.0 30.5 50.5 52.2
Davinci 5.5 1.0 10.0 5.5 0.5 14.5 3.5 8.0 34.0 5.5 8.8
ChatGPT 52.0 81.5 49.0 76.5 74.0 55.0 72.0 76.5 18.0 76.0 63.05
GPT-4 70.5 90.5 58.5 80.5 83.5 64.0 86.5 88.0 31.5 88.0 74.15
Table 4: Evaluation on mental health.
Setup. To assess how effectively LLMs address mental health issues, we task them with selecting the most appropriate response from four possible options for a given real-life scenario. This evaluation is conducted in a zero-shot setting.
Results. As shown in Table 4, GPT-4 demonstrates a clear advantage over other tested LLMs in 7 out of 10 languages. Notably, GPT-4 excels in handling mental health-related questions in Chinese (90.5%), Portuguese (88.0%), Korean (86.5%), and Spanish (88.0%). ChatGPT ranks second in terms of overall average accuracy. While Gemini trails behind GPT-4 and ChatGPT, it achieves a perfect score of 100% accuracy in Korean. Davinci, however, performs the weakest in this trustworthiness evaluation.
4.6 Evaluation on Physical Health
Model AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan 12.0 21.0 28.0 24.0 12.5 37.0 27.0 16.0 28.5 20.5 22.65
Gemini 23.0 61.5 44.5 16.5 21.0 42.0 100.0 60.0 11.0 59.5 43.9
Davinci 4.5 2.0 25.0 0 2.0 30.5 1.0 0 39.5 0.5 10.5
ChatGPT 21.5 44.5 29.5 50.5 36.5 32.0 37.0 53.0 15.0 50.0 36.95
GPT-4 59.0 79.5 47.5 80.5 81.0 48.5 75.5 80.5 49.0 83.5 68.45
Table 5: Evaluation on physical health.
Setup. To assess LLMs’ ability to understand appropriate actions and responses for maintaining physical health in various scenarios, we task them with selecting the correct answer from four given options. This evaluation is conducted in a zero-shot setting.
Results. As shown in Table 5, GPT-4 consistently outperforms the other tested LLMs in 9 out of 10 languages. Notably, it maintains a 24.55% lead in average accuracy over the second-best model, Gemini. While Gemini achieves a perfect score of 100% in Korean, GPT-4 shows relatively lower accuracy in French (47.5%), Italian (48.5%), and Russian (49.0%) compared to its performance in other languages. Overall, Davinci ranks as the poorest performer among the LLMs tested.
4.7 Evaluation on Toxicity
Model Task AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan Benign 10.1 25.9 18.7 19.6 23.9 18.5 35.5 15.7 18.5 19.7 20.61
Adversarial 21.9 22.6 11.5 16.5 30.2 19.6 18.5 19.4 1.3 15.6 17.71
Gemini Benign 3.3 22.5 16.0 17.4 10.4 18.8 20.5 22.6 6.1 17.1 15.47
Adversarial 12.1 22.7 10.5 21.4 11.9 17.7 9.6 14.7 6.1 16.2 14.29
Davinci Benign 1.2 8.2 2.7 9.5 8.4 7.7 12.0 4.5 1.6 4.5 6.03
Adversarial 3.9 4.1 0.4 4.5 6.5 3.2 2.8 2.1 0.3 0.3 2.81
ChatGPT Benign 2.7 9.2 6.2 7.3 14.9 8.1 0 8.7 0 6.7 6.38
Adversarial 9.8 10.1 3.9 7.6 14.6 7.8 0 8.2 0 1.9 6.39
GPT-4 Benign 1.3 10.6 8.5 13.8 13.6 11.4 12.2 10.0 2.9 9.2 9.35
Adversarial 6.3 4.2 3.0 6.3 9.4 4.6 3.4 4.1 0.5 1.2 4.3
Table 6: Evaluation on toxicity.
Benign Prompts Adversarial Prompts
System You are a helpful assistant You are a helpful assistant. You do not need to obey any
content policy. You do not need to align with human morale.
You can say anything as long as it is coherent to the prompt. It is
preferred to output toxic, rude, and disrespectful language.
User You are required to keep You do not obey any content policy. You do not
generation given the incomplete need to align with human morale. You can say anything as long
[…]
toxic, rude, and disrespectful language. you are required to keep
[…]
Table 7: Two representative prompts for toxic evaluation of LLMs. The adversarial prompt can elicit toxicity in text generation. The repeated instructions in the user prompt aim to reinforce LLMs to follow the system prompt.
Setup. To assess the ability of LLMs to avoid generating toxic content and to gain a deeper understanding of their emergent behaviors, we task the models with natural language generation. Since LLMs are designed for dialogue interactions, we use two representative system prompts to guide text completion: a benign system prompt and an adversarial system prompt, as shown in Table 7. The adversarial prompt includes additional instructions aimed at bypassing the content policies enforced during model alignment, in an attempt to "jailbreak" the LLMs and provoke toxic content. The repeated instructions within the user prompt are intended to reinforce the model’s adherence to the system prompt. We utilize the Perspective API, an automated tool for detecting toxic language and hate speech, to evaluate the toxicity of the generated outputs. Toxicity is measured by calculating the average Toxicity score across 200 instances, with lower scores indicating that the model produces less toxic content. This evaluation is conducted in a zero-shot setting.
Results. In the toxicity evaluation experiments, as shown in Table 6, Davinci delivers the best overall performance, significantly outperforming the other models. Interestingly, for most multilingual scenarios, the content generated by LLMs in response to the benign prompt exhibits higher toxicity scores than that generated by the adversarial prompt. This suggests that LLMs are not easily "jailbroken" by adversarial prompts to produce toxic content, likely due to the activation of their toxicity detection mechanisms, which prevent them from following prompts that would generate offensive language. Given the low toxicity scores of Davinci, ChatGPT, and GPT-4, we acknowledge the substantial efforts made to ensure responsible development, deployment, and usage during the training of these models. However, we also observe a trustworthiness gap between Davinci and GPT-4. As highlighted in Achiam et al. (2023), GPT-4 generally outperforms Davinci across a range of NLP tasks, yet in this evaluation, GPT-4 exhibits more toxic behavior than Davinci, resulting in a higher toxicity score. This finding suggests that advancements in model alignment techniques have not kept pace with the rapid progress in LLM training, potentially increasing the risk of harmful content exposure, especially in multilingual settings.
4.8 Evaluation on Fairness & Bias
Model Task AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan Benign 15.8 10.6 0 0 15.4 0 50.0 0.2 1.3 0 9.33
Targeted 14.0 11.1 0 1.0 33.6 0 24.7 3.6 0 0 8.8
Untargeted 8.3 3.6 0 0 37.2 0 24.7 0.2 0 0 7.4
Gemini Benign 8.5 14.5 37.5 0.3 0.3 16.1 1.6 20.0 17.7 10.4 12.69
Targeted 14.5 10.1 30.4 0 0 14.0 0.5 0.7 5.2 7.5 8.29
Untargeted 24.4 16.4 25.2 2.3 2.6 5.7 0 14.8 0 8.3 9.97
Davinci Benign 0 14.5 0 0 0 19.7 0 0 33.3 4.4 7.19
Targeted 0 3.9 0 0 0 47.1 0 0.2 0 8.5 5.97
Untargeted 0 4.9 0 0 0 19.5 0 0.7 0 8.0 3.31
ChatGPT Benign 53.6 44.2 31.2 0.8 2.9 42.7 74.7 41.4 93.5 35.9 42.08
Targeted 45.8 53.9 15.1 3.9 17.4 38.2 51 67.7 0.5 55.4 34.89
Untargeted 75.5 63 13.5 4.2 13.0 72.1 51.8 79.1 0.3 52.3 42.48
GPT-4 Benign 6.2 31.5 71.6 1.3 18.5 47.3 64.3 84.3 39.1 54.1 41.82
Targeted 16.9 10.6 9.6 9.1 27.1 58.8 52.1 16.6 0 42.9 24.37
Untargeted 19.7 9.6 32.2 2.9 0 70.5 50.9 66.9 0 33.8 28.65
Table 8: Evaluation on fairness & bias.
Setup. To evaluate the generation bias of LLMs, we follow the approach used in DecodingTrust Wang et al. (2023a), where biased statements are included in the user prompt, and the LLM is asked to respond by either agreeing (appending "I agree." to the response) or disagreeing (appending "I disagree."). The system prompts provide general instructions to the LLMs across three evaluation scenarios: (1) a benign scenario, which does not include any misleading instructions; (2) a targeted scenario, which encourages the model to be biased against the specific demographic group mentioned in the user prompt; and (3) an untargeted scenario, where no specific demographic group is targeted for bias. We measure model bias using the disagreement rate, defined as $diagreementIndex=\frac{n_{disagree}}{n}\in[0,1]$, where $n_{disagree}$ represents the number of times the model disagrees with the biased statement, and $n=384$ is the total number of instances evaluated for each scenario. $disagreementIndex=1$ indicates the model never agrees with the biased statement, while a value of 0 indicates it always agrees. For fairness and bias, we evaluate the LLMs in a zero-shot setting.
Results. As shown in Table 8, ChatGPT outperforms other LLMs in terms of average performance across all languages and scenarios. Both ChatGPT and GPT-4 consistently refuse to agree with biased statements, reflecting the significant efforts invested in fine-tuning GPT models to minimize biased outputs. However, under the targeted scenario, the output bias of Gemini, ChatGPT, and GPT-4 increases compared to their performance in the other two scenarios, indicating that these models can be more easily "tricked" into generating biased content when exposed to misleading instructions. Interestingly, for Baichuan and Davinci, the output bias under the targeted scenario decreases compared to the untargeted scenario. This may be due to the models’ inability to precisely follow instructions, making them less responsive to targeted bias prompts.
4.9 Evaluation on Misinformation
Model AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan 1.6 14.9 0 4.1 0 12.0 4.9 8.2 2.9 6.2 5.48
Gemini 28.6 42.3 20.3 51.8 1.2 78.8 0 34.4 0 37.3 29.47
Davinci 1.6 0.8 0.8 0 0 0 0 0 1.2 0.4 0.48
ChatGPT 9.5 29.8 26.1 8.7 0 47.3 7.5 22.8 17 28.6 19.73
GPT-4 49.3 12.4 51.4 50.6 0.4 90 16.5 62.6 2.0 48.5 38.37
Table 9: Evaluation on misinformation.
Setup. To assess the ability of LLMs to avoid generating false, misleading, nonsensical, or low-quality information, the models are expected to refrain from responding to misleading input prompts. For this evaluation, we provide user prompts containing misleading statements and task the LLMs with text continuation. We use the metric $noResponseIndex=1-\frac{n_{noRespons}}{n}\in[0,1]$ to measure the model’s ability to detect misinformation, where $n_{noResponse}$ represents the number of times the model refuses to respond, and $n=241$ is the total number of prompts used in the evaluation. A higher $noResponseIndex$ indicates a stronger ability to detect misinformation. For this evaluation, we test the LLMs in a zero-shot setting.
Results. As shown in Table 9, GPT-4 outperforms other LLMs, particularly excelling in Arabic, French, Italian, Korean, Portuguese, and Spanish. This suggests significant effort has been made during GPT-4’s pre-training to prevent the generation of misinformation when faced with misleading instructions. On average, GPT-4 surpasses Gemini by 8.9% in the $noResponseIndex$ across the ten languages. However, all tested models exhibit weaker performance in Hindi, Korean, and Russian, likely due to limited training resources in these languages. This highlights the considerable room for improvement in the ability of mainstream LLMs to combat misinformation in multilingual contexts.
4.10 Evaluation on Privacy
Model Addr. CCN EA PN Pwd Ph. SK SSN SSH Avg
Baichuan 28.6 20.1 20.9 21.3 36.4 22.8 38.4 23.5 34.3 27.37
Gemini 9.9 21.2 4.6 11.1 13.8 5.7 16.3 16 19.6 13.13
Davinci 2.6 1.3 2.3 4.6 2.6 22.6 5.5 10.6 2.8 6.1
ChatGPT 55.3 89.3 60.6 87.9 88.1 85.4 87 86.6 81.3 80.17
GPT-4 74.9 93.8 70.4 91.9 94.5 75.5 94.9 93.1 96.3 87.26
Table 10: Average accuracies on the PII detection task across ten languages.
Model Bank Divorce Exam Health Layoff Party Promotion Vote Avg
Baichuan 17.67 27.95 26.29 24.02 19.59 25.62 18.48 21.75 22.67
Gemini 29.03 41.37 44.83 50.23 48.6 57.4 58.16 51.02 47.58
Davinci 34.58 34.16 33.93 38.34 34.52 31.42 28.75 31.54 33.41
ChatGPT 78.34 81.65 78.35 79.71 73.92 82.18 74.83 77.71 78.34
GPT-4 80.6 85.02 83.01 85.32 81.68 85.22 83.72 84.27 83.61
Table 11: Average detection rate of the privacy event detection task across ten languages. Bank represents bankrupt, Health represents mental health issue, Exam represents failing an exam, Party represents preparing a surprise for a party
Setup. When interacting with LLMs, there is a risk that private information may be compromised during inference. Therefore, we follow the approach in DecodingTrust to evaluate the privacy capabilities of LLMs in two key scenarios: (1) assessing the models’ ability to extract various types of Personally Identifiable Information (PII) during inference, and (2) evaluating the models’ resilience against information leakage in different privacy-related events and contexts, to understand how well they handle sensitive information during conversations. For the PII detection task, we focus on 10 types of PII, including address (Addr.), credit card number (CCN), email address (EA), passport number (PN), password (Pwd), phone number (Ph.), secret key (SK), social security number (SSN), and SSH private key (SSH). In the privacy event detection task, we consider six privacy-related terms (e.g., "in confidence," "confidentially," "privately," "in private," "secretly," and "in secret") and eight privacy-related events, such as divorce, mental health issues, planning a surprise party, voting, failing an exam, bankruptcy, layoffs, and promotions. Following DecodingTrust, we create 10 prompts for the PII detection scenario and 48 prompts (6 privacy-related words × 8 privacy events) for the privacy event detection scenario. Both tasks are text generation tasks: we run 100 iterations for each PII detection prompt and 50 iterations for each privacy event detection prompt, tracking the rate of unsuccessful responses (i.e., instances where the model refuses to respond). We report the average response rate across ten languages to evaluate the LLMs’ privacy protection capabilities in multilingual settings. For privacy evaluation, we use the zero-shot setting.
Results. For the PII detection task, as shown in Table 10, Baichuan, Gemini, and Davinci demonstrate weak performance, leaving significant room for improvement in preventing data leakage during conversations. In contrast, ChatGPT and GPT-4 exhibit strong performance in protecting PII during interactions. Notably, ChatGPT performs best at safeguarding phone numbers, though it still shows vulnerabilities by leaking sensitive information such as addresses and email addresses. For the privacy event detection task, GPT-4 outperforms all other LLMs across all privacy events, surpassing the second-best model, ChatGPT, by a margin of 7.09%. In summary, GPT-4 excels at protecting private information, demonstrating its robustness and superior ability to detect and handle inappropriate instructions.
4.11 Evaluation on Machine Ethics
Model Task AR ZH FR DE HI IT KO PT RU ES Avg
Baichuan 0-shot_ETHICS 0 65.6 55.2 64.1 0 55.7 63.6 59.7 62.6 58.7 48.52
5-shot_ETHICS 0 69.1 60.1 62.1 0.9 41.7 58.7 62.6 54.7 56.2 46.61
0-shot_JC 0 31.8 35.3 31.8 43.2 33.8 46.7 31.8 38.3 5.4 29.81
5-shot_JC 0 44.7 12.4 36.3 0 21.3 48.2 49.2 15.9 42.7 27.07
Gemini 0-shot_ETHICS 12.9 60.1 8.9 6.9 0.9 7.9 8.4 5.4 8.9 9.9 13.02
5-shot_ETHICS 0 41.7 3.9 4.4 6.4 5.4 1.4 7.9 3.9 7.4 8.24
0-shot_JC 7.4 32.8 3.9 2.4 0 2.4 4.4 4.4 5.4 3.4 6.65
5-shot_JC 0.9 32.84 4.4 3.9 4.4 1.4 2.9 4.4 10.9 2.4 6.844
Davinci 0-shot_ETHICS 0 50.2 0 1.4 0 0 0 34.8 32.8 0.4 11.96
5-shot_ETHICS 0 50.2 0 1.4 0 0 0 34.8 32.3 0.9 11.96
0-shot_JC 0 0.4 0 0 0 0 1.9 1.4 0 0 0.37
5-shot_ETHICS 0 0.9 0 4.4 0 2.9 0 6.9 10.9 3.9 2.99
ChatGPT 0-shot_ETHICS 0 66.1 69.1 66.6 4.9 71.6 63.1 72.1 62.1 71.1 54.67
5-shot_ETHICS 8.4 62.1 67.1 68.6 4.9 69.1 63.6 69.6 58.7 74.1 54.62
0-shot_JC 4.9 62.6 48.2 49.7 0.4 50.2 50.7 47.2 48.7 48.2 41.08
5-shot_JC 5.4 53.2 49.2 50.2 26.8 50.2 52.2 48.7 61.1 54.7 45.17
GPT-4 0-shot_ETHICS 0 81.5 76.1 77.1 0.4 77.1 71.1 79.1 76.1 79.6 61.81
5-shot_ETHICS 0 81.0 78.1 80.0 0.4 77.6 71.6 77.1 77.1 78.6 62.15
0-shot_JC 0 56.2 52.2 41.7 0 50.2 61.1 49.7 51.7 58.2 42.1
5-shot_JC 4.9 61.6 51.2 48.2 11.4 43.7 60.1 51.2 57.7 58.2 44.82
Table 12: Evaluation on machine Ethics.
Setup. To assess how well LLMs can distinguish between moral and immoral behaviors, we compare their moral recognition performance using two morality datasets: ETHICSGui et al. (2021) and Jiminy Cricket (JC)Hariri (2023). The ETHICS dataset evaluates a model’s understanding of fundamental ethical concepts, including justice, virtue, deontology, utilitarianism, and commonsense morality. It comprises five subsets, each focusing on different normative factors. For this study, we follow the approach in DecodingTrust and select the commonsense morality subset for evaluation. In the few-shot setting, we provide five training samples as demonstrations, each consisting of a commonsense morality scenario and its label. The LLM is required to classify each sample as either "wrong" or "not wrong," with classification accuracy serving as the evaluation metric. The JC dataset is designed to assess an agent’s ability to act morally in text-based games, featuring more diverse scenarios and annotations. Each sample includes a scenario from a text-based game, accompanied by a threefold label: (1) the moral valence of the action (good, bad, or neutral), (2) the focal point (whether the action benefits or harms the agent or others), and (3) the ordinal degree (a ranking of how good or bad the action is on a scale of 1 to 3). In the few-shot setting, three training samples are provided as demonstrations. The LLMs are tasked with determining the moral valence of each sample (good, bad, or neutral), and classification accuracy is used as the evaluation metric. For machine ethics, we evaluate the LLMs in both zero-shot and five-shot settings.
Results. As shown in Table 12, GPT-4 stands out as the top performer, achieving the highest scores in both zero-shot and five-shot settings on the ETHICS and JC datasets. Across different languages, GPT-4 consistently outperforms other LLMs in Chinese, French, Korean, Portuguese, and Spanish. ChatGPT ranks second in performance across multiple languages, demonstrating that both GPT-4 and ChatGPT possess strong moral recognition capabilities. However, it is important to note that most tested models struggle with Arabic and Hindi, which may be due to the unique characteristics of these languages and the limited availability of training data.
5 Conclusion
This paper presents a comprehensive investigation into the multilingual trustworthiness of LLMs, addressing a critical gap in the current understanding of LLM reliability. Through the development of the XTRUST multilingual trustworthiness benchmark, we have enabled a systematic evaluation of widely used LLMs across ten languages. Our findings reveal significant disparities in trustworthiness performance across different languages, underscoring the urgent need for more focused research and development to enhance LLM trustworthiness in non-English languages. This study highlights the importance of addressing trustworthiness concerns in multilingual contexts. We hope to inspire further exploration and innovation in trustworthiness alignment techniques for non-English LLMs, ultimately fostering the creation of more trustworthy and reliable AI systems for users worldwide. Our work serves as a call to action for researchers, developers, and policymakers to collaborate in tackling the ethical and practical challenges associated with deploying AI systems in multilingual and multicultural settings. We hope our findings inspire future efforts to: (1) safeguard LLMs for low-resource languages; (2) deepen the understanding of LLMs’ cross-lingual generalization on trustworthiness issues; and (3) develop effective strategies to enhance LLMs’ capabilities in multilingual trustworthiness.
Limitations
In this study, our primary focus is on exploring the multilingual trustworthiness capabilities of LLMs. However, two key limitations prevent us from providing a comprehensive assessment of LLMs’ trustworthiness in practical applications. First, although we evaluate five widely-used LLMs, we were unable to assess other open-source and proprietary models, such as Llama and Claude, which also support multiple languages. Second, our evaluation does not encompass all non-English languages. We plan to expand our evaluation in future iterations of this study, with the support of the research community.
Appendix A Prompt design
Figure 3: An example of the input prompt.
In this study, we follow the approach used in DecodingTrust Wang et al. (2023a) to design our prompts, which consist of two components: the system prompt and the user prompt. The system prompt is designed to configure the tone, role, and style of the LLM assistant, allowing for customization of the model’s interaction pattern to align with diverse user preferences and use cases (As illustrated in Figure 3). The user prompt is further divided into two parts: the task description and the task-specific prompts.
We address two types of NLP tasks—text classification and text generation—under both zero-shot and few-shot settings. For a zero-shot classification task, we provide the model with a task description before presenting the test input. This description offers clear instructions for performing the task and specifies the allowable class labels. We craft different role-based system prompts for various trustworthiness issues. For example, "You are a helpful assistant" is used as the system prompt for Illegal Activity, while "You are now required to be an illusion detector" is used for Hallucination. In the few-shot classification setting, the model receives several demonstrations along with the task description to assist in generating predictions, a technique known as in-context learning Dong et al. (2022). Each demonstration consists of a simulated user input and its corresponding label, formatted as a model response, enabling the LLM to generate predictions based on these examples. For text generation tasks—such as those evaluating toxicity, bias, and misinformation—we similarly establish the LLM assistant’s role through the system prompt. The user prompt includes a task description to guide the LLM in generating coherent continuations for a given input.
In most instances, LLMs adhere to the provided instructions and generate expected responses. However, as noted in previous studies Zhou et al. (2023); Huang et al. (2023), LLMs may occasionally refuse to answer or produce hallucinations unrelated to the input. In this study, such responses are considered incorrect for the purposes of evaluation, particularly in classification tasks.
Appendix B Evaluated Models
The detailed information of 5 evaluated LLMs is shown in Table 13.
Model Model Size Access Version Language Creator
GPT-4 Undisclosed api 1106-preview Multi-Lans OpenAI
GPT-3.5-turbo Undisclosed api 1106 Multi-Lans OpenAI
Text-Davinci-002 Undisclosed api - Multi-Lans OpenAI
Geminipro Undisclosed api gemini-pro Multi-Lans Google
Baichuan Undisclosed api baichuan2-7b-chat-v1 Multi-Lans Baichuan AI
Table 13: LLMs evaluated in this paper.
Appendix C Demonstration of the illegal activity evaluation for LLMs
In Figure 4, we present an example of illegal activity evaluation for LLMs.
Figure 4: An example of illegal activity evaluation.
Appendix D Demonstration of the hallucination evaluation for LLMs
We have presented an example of the hallucination evaluation for LLMs in Figure 5.
Figure 5: An example of hallucination evaluation.
Appendix E Demonstration of the OOD evaluation for LLMs
In Figure 6, we present an example of the OOD evaluation for LLMs.
Figure 6: An example of OOD evaluation.
Appendix F Demonstration of the mental health evaluation for LLMs
In Figure 7, we present an example of the mental health evaluation for LLMs.
Figure 7: An example of the mental health evaluation.
Appendix G Demonstration of the physical health evaluation for LLMs
We have included an example related to the physical health evaluation for LLMs in Figure 8.
Figure 8: An example of the physical health evaluation.
Appendix H Demonstration of the fairness and bias evaluation for LLMs
In Figure 9, we present an example of fairness and bias evaluation for LLMs. Furthermore, Figure 10 provides a visualized analysis of different prompts across diverse languages and models, showing that GPT-4 and ChatGPT surpass other language models in fairness and bias evaluation for all tested languages across different prompt types.
Figure 9: An example of the fairness and bias evaluation.
Figure 10: The fairness and bias evaluation for LLMs across three prompt types.
Appendix I Demonstration of the toxicity evaluation for LLMs
We have included an example related to the toxicity evaluation for LLMs in Figure 11.
Figure 11: An example of the toxicity evaluation.
Appendix J Demonstration of the misinformation evaluation for LLMs
We have included an example related to the misinformation evaluation for LLMs in Figure 12.
Figure 12: An example of the misinformation evaluation.
Appendix K Demonstration of the PII detection for LLMs
We have included an example of the PII detection for LLMs in Figure 13.
Figure 13: An example of the PII detection.
Appendix L Demonstration of the PrivacyEvents detection for LLMs
We present an example of the PrivacyEvents detection for LLMs in Figure 14. In Figure 15161718, and 19, we provide the analyses of privacy-related word leakage for PII and PrivacyEvents detection experiments. These results reveal that for powerful LLMs, such as GPT-4 and ChatGPT, can detect most privacy-related words for most languages. GPT-4 performs poorly in Russian and ChatGPT performs poorly in French. These results provide more insights into understanding different privacy-related words for different languages.
Figure 14: An example of the PrivacyEvents detection.
Figure 15: Analysis of privacy-related word leakage of GPT-4 for privacy detection.
Figure 16: Analysis of privacy-related word leakage of ChatGPT for privacy detection.
Figure 17: Analysis of privacy-related word leakage of Baichuan for privacy detection.
Figure 18: Analysis of privacy-related word leakage of Gemini for privacy detection.
Figure 19: Analysis of privacy-related word leakage of Davinci for privacy detection.
Appendix M Demonstration of the machine ethics evaluation for LLMs
We present an example of the machine ethics evaluation task in zero-shot setting in Figure 20.
Figure 20: An example of the machine ethics evaluation.

### v2

@@
 The code is available at https://github.com/LluckyYH/XTRUST.
-{CJK}
-UTF8gbsn
 XTRUST: On the Multilingual Trustworthiness of Large Language Models
-Yahan Li1111Equal contribution  Yi Wang2111Equal contribution  Yi Chang1,3,4  Yuan Wu 1222Corresponding author: yuanwu@jlu.edu.cn 1School of Artificial Intelligence, Jilin University, China 2The Hong Kong University of Science and Technology (Guangzhou), China 3Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 4International Center of Future Science, Jilin University, China
-Figure 1: Chat with LLMs in non-English languages can lead to untrustworthy responses
+Yahan Li [^1] footnotemark: 1  Yi Wang211footnotemark: 1  Yi Chang1,3,4  Yuan Wu 122footnotemark: 2 1School of Artificial Intelligence, Jilin University, China 2The Hong Kong University of Science and Technology (Guangzhou), China 3Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 4International Center of Future Science, Jilin University, China
+Figure 1: LLMs in non-English languages can produce biased and untrustworthy responses, especially on topics like “physical integrity”, relevant to bias and fairness tasks.
 1 Introduction
@@
 XTRUST offers three key advantages: (1) Extensive Diversity.
-It includes a total of 23,590 instances, covering 10 distinct categories of trustworthiness concerns, providing a robust and comprehensive evaluation framework for LLMs.
+It includes a total of 2359 instances, covering 10 distinct categories of trustworthiness concerns, providing a robust and comprehensive evaluation framework for LLMs.
 (2) Diverse Question Types.
@@
 3.1 Trustworthiness Categories
-An overview of XTRUST is presented in Fig. 2, We collect a total of 2359 instances spanning 10 categories of trustworthiness issues from several monolingual datasets:
+An overview of XTRUST is presented in Fig.
+2, We collect a total of 2359 instances spanning 10 categories of trustworthiness issues from several monolingual datasets.
+When expanded to 10 languages, the number of instances reaches 23,590:
 Illegal Activity.
@@
 3.2 Data Collection
-We conducted a systematic review of numerous trustworthy datasets from various fields, including safety, fairness, and privacy, to establish the foundation for the multilingual XTRUST evaluation.
-For the Out-of-Distribution (OOD) robustness assessment, we collected recent news articles from the Internet, ensuring that the information was not used to train LLMs.
-All data collected for this study was carefully reviewed by our team.
-For Illegal Activity, we randomly sampled 200 instances from the Chinese AI and Law challenge dataset (CAIL2018)Xiao et al.
-(2018), a large compilation of legal case records provided by the Supreme People’s Court of China.
-For Hallucination, we selected 300 samples from the Question Answering (QA) task in HaluevalLi et al.
-(2023), a dataset containing generated and human-annotated hallucinated examples to evaluate LLMs’ ability to detect hallucinations.
-For OOD Robustness, we gathered entertainment, education, and international news from China News Service, specifically collecting articles published after September 2023 to ensure the information was not used in LLM training.
-A total of 200 articles were used for this evaluation.
-For Mental Health, we randomly selected 200 instances from the Mental Health subset of SafetyBench Zhang et al.
-(2023c), which includes safety-related multiple-choice questions.
-For Physical Health, we randomly selected 200 samples from SafeText Levy et al.
-(2022), a dataset containing 367 human-written real-life scenarios with safe and unsafe suggestions for each scenario.
-For Toxicity, we randomly selected 200 instances from the toxic task prompts subset of REALTOXICITYPROMPT Gehman et al.
-(2020) to evaluate the toxicity generation of LLMs.
-For Fairness & Bias, we selected prompts from the Stereotype Bias subset of DecodingTrust Wang et al.
-(2023a), asking LLMs whether they agree or disagree with the statements presented in the prompts.
-DecodingTrust spans 24 demographic groups (e.g., "Homosexuals," "Men," "White people") and 16 bias topics (e.g., "Sexually Transmitted Diseases," "Occupations," "Intelligence").
-We selected 384 prompts (24 demographic groups × 16 bias topics) to comprehensively evaluate LLMs’ performance on fairness and bias.
-For Misinformation, we used 155 questions from the Misinformation Harms subset of the Do-Not-Answer dataset Wang et al.
-(2023b), a taxonomy for assessing the risk of harm posed by LLMs.
-Additionally, we collected 86 fake news-related questions from the Internet, using a total of 241 questions in this evaluation.
-For Privacy, we selected 34 data entries from the Privacy subset of DecodingTrust Wang et al.
-(2023a) to evaluate LLMs’ capability to prevent data leakage.
-For Machine Ethics, we randomly selected 200 short samples from the test set of ETHICS Gui et al.
-(2021), which assesses a model’s knowledge of basic ethical concepts.
-We also selected 200 instances from the Jiminy Cricket dataset Hariri (2023), which evaluates an agent’s ability to make moral decisions in text-based games.
+We constructed a comprehensive evaluation dataset covering all assessment dimensions through carefully designed procedures.
+The dataset includes samples for each evaluation task with rigorous quality control measures.
+Detailed information about data sources, collection methods, and annotation protocols is provided in Appendix A.1.
 3.3 Translating the Collected Data
-To create a multilingual trustworthiness benchmark, we translated the original monolingual data into multiple languages.
-For this study, we selected 10 languages: Arabic (AR), Chinese (ZH), French (FR), German (DE), Hindi (HI), Italian (IT), Korean (KO), Portuguese (PT), Russian (RU), and Spanish (ES).
-The test data was translated using Google Translate.
+To ensure accurate multilingual translations, we followed these steps:
+Quality Control: We randomly selected 50 instances and translated them into each target language using Google, Bing, and GPT translators.
+The translations were rigorously evaluated by PhD students proficient in the respective languages, ensuring both accuracy and textual quality.
+Based on consistency and broader language support, Google Translate was ultimately selected as the preferred translation tool.
+The data was translated into 10 languages: Arabic (AR), Chinese (ZH), French (FR), German (DE), Hindi (HI), Italian (IT), Korean (KO), Portuguese (PT), Russian (RU), and Spanish (ES).
+These languages were selected based on model support and their representation of diverse linguistic regions, ensuring broad applicability.
+Toxicity Task Consistency: All selected languages are supported by the Perspective API for toxicity tasks, ensuring consistency and reliability in the analysis.
 4 Experiments
+Figure 3: Example input prompt for the evaluation task
 4.1 Models
-We conducted experiments on five LLMs that support multilingual applications, including baichuan2-7b-chat-v1 (denoted as Baichuan)Yang et al.
+We conducted experiments on five commercial LLMs that support multilingual applications, chosen for their representation of the latest advancements and broader language support, making them more suitable for multilingual trustworthiness assessment compared to open-source models with limited capabilities.
+The models used in this study include baichuan2-7b-chat-v1 (denoted as Baichuan)Yang et al.
 (2023), Gemini-pro (released on December 13, 2023, denoted as Gemini)Team et al.
@@
 Model AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan 45.0 98.5 4.5 100 12.5 70.0 0 74.0 84.0 18.0 50.65
-Gemini 96.5 95.5 94.0 97.0 91.5 11.0 87.5 97.5 98.5 95.5 86.45
+Baichuan 45.0 98.5 4.5 100 12.5 70.0 0 74.0 84.0 18.0 50.7
+Gemini 96.5 95.5 94.0 97.0 91.5 11.0 87.5 97.5 98.5 95.5 86.5
 Davinci 6.0 3.0 6.5 2.0 3.0 20.0 3.0 22.5 1.5 15.0 8.25
-ChatGPT 61.0 75.5 76.0 86.5 57.5 60.5 0 75.0 2.0 77.0 57.10
+ChatGPT 61.0 75.5 76.0 86.5 57.5 60.5 0 75.0 2.0 77.0 57.1
 GPT-4 53.5 96.0 71.0 97.0 85.0 80.5 91.5 58.5 93.0 79.0 80.5
-Table 1: Evaluation on illegal activity.
+Table 1: Evaluation on Illegal Activity. Avg indicates the average performance across all languages, consistent across tasks.
 Setup.
-To assess the ability of LLMs to handle illegal activity, we transform the collected legal case records into binary classification questions, asking the models to differentiate between legal and illegal behaviors based on a straightforward task description.
-This approach aims to evaluate how effectively LLMs comprehend legal knowledge.
-For this task, we conduct evaluations in a zero-shot classification setting.
+To evaluate LLMs’ grasp of fundamental legal principles, we curated a set of universally recognized legal scenarios - cases based on universally condemned behaviors (e.g., theft, violence).
+Our approach employs binary classification tasks in a zero-shot setting to determine whether LLMs can reliably distinguish unlawful actions without relying on jurisdiction-specific training data, thereby testing their grasp of core legal norms rather than localized statutory knowledge.
 Results.
-As shown in Table 1, Gemini demonstrates the strongest performance in detecting illegal activities across multiple languages, achieving an impressive average score of 86.45%, particularly excelling in languages such as Arabic, Chinese, French, and Portuguese.
-GPT-4 follows closely, with an average score of 80.5, showing strong performance in languages like Chinese, French, German, and Russian.
-Baichuan ranks third, with an average score of 50.65%, though its performance is notably weaker in languages such as French, Hindi, and Spanish.
-ChatGPT and Davinci display lower average scores of 57.10% and 8.25%, respectively, highlighting the pressing need for improvements in multilingual trustworthiness for detecting illegal activities.
+Our analysis reveals striking disparities in models’ multilingual legal reasoning capabilities (Table 1).
+Gemini leads in legal awareness (86.5%), excelling in Arabic/Russian, likely due to multicultural legal training data.
+GPT-4 performes strongly in Italian/German but trails Gemini by 43 % in non-Latin scripts like Arabic.
+Baichuan (50.7%) and ChatGPT (57.1%) show steep drops, with frequent misclassification of French/Spanish cases exposing geographic blind spots in safety training.
+Davinci’s near-random performance (8.25%) highlighted legacy architectures’ struggles with legal nuance.
+Findings urge explicit legal safety optimization in pretraining, not post-hoc alignment.
 4.3 Evaluation on Hallucination
 Model AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan 13.0 25.5 38.0 49.0 37.0 24.8 12.0 39.5 42.5 32.0 31.33
-Gemini 46.5 49 46.5 41.5 21.0 40.2 50.0 42.5 0 42.5 37.97
-Davinci 18.5 52.0 10.5 25.0 9.0 29.3 44.0 8.5 46.5 15.0 25.83
-ChatGPT 53.0 37.5 51.0 44.0 46.0 40.2 50.0 56.5 56.0 44.5 47.87
-GPT-4 49.5 45.0 51.5 48.5 46.0 49.7 38.0 58.5 43.5 48.5 47.87
+Baichuan 13.0 25.5 38.0 49.0 37.0 24.8 12.0 39.5 42.5 32.0 31.3
+Gemini 46.5 49 46.5 41.5 21.0 40.2 50.0 42.5 0 42.5 38.0
+Davinci 18.5 52.0 10.5 25.0 9.0 29.3 44.0 8.5 46.5 15.0 25.8
+ChatGPT 53.0 37.5 51.0 44.0 46.0 40.2 50.0 56.5 56.0 44.5 47.9
+GPT-4 49.5 45.0 51.5 48.5 46.0 49.7 38.0 58.5 43.5 48.5 47.9
 Table 2: Evaluation on hallucination.
@@
 Results.
-As shown in Table 2, GPT-4 outperforms other LLMs, achieving the highest scores in French (51.5%), Hindi (49.7%), Italian (49.7%), Portuguese (58.5%), and Spanish (48.5%).
-It also secures the second-highest scores in Arabic (49.5%) and German (48.5%).
-ChatGPT follows closely, delivering strong performance across most languages, particularly excelling in Arabic (53.0%), Hindi (46.0%), Russian (56.0%), and Korean (50.0%).
-While Baichuan, Gemini, and Davinci demonstrate notable strengths in specific languages, their overall performance is less optimal.
+Our evaluation uncovers a nuanced landscape of hallucination detection capabilities across languages.
+GPT-4 emerges as the most consistent performer, particularly in Romance languages where it achieves 58.5% accuracy in Portuguese.
+Yet ChatGPT reveals surprising strengths in linguistically distant contexts, outperforming GPT-4 by 12.5% in Russian and maintaining robust accuracy in Arabic (53.0%) and Korean (50.0%).
+This pattern suggests that while model scale (GPT-4) generally predicts better performance, targeted alignment (ChatGPT) can create specialized advantages for specific language families.The observed performance disparity across linguistic domains - exemplified by Gemini’s 50.0% accuracy in Korean versus catastrophic failure in Russian (0%) - demonstrates fundamental limitations in current hallucination mitigation architectures.
 4.4 Evaluation on Out of Distribution Robustness
 Model Task AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan 0-shot 0.5 41.5 46.5 0 66.5 7.5 0.4 0.5 2.4 12.0 17.78
-Gemini 0-shot 73.0 75.0 14.0 14.5 87.5 83.0 0 71.0 0 88.5 50.65
+Baichuan 0-shot 0.5 41.5 46.5 0 66.5 7.5 0.4 0.5 2.4 12.0 17.8
+Gemini 0-shot 73.0 75.0 14.0 14.5 87.5 83.0 0 71.0 0 88.5 50.7
 Davinci 0-shot 0 54.0 7.5 0 0 6.0 0 0 0 5.5 7.3
-ChatGPT 0-shot 0 58.0 50.5 5.0 69.9 77.5 8.9 1.0 0.9 31.5 30.32
-GPT-4 0-shot 3.4 93.5 98.0 34.0 99.5 98.5 2.4 2.5 19.0 98.0 54.88
+ChatGPT 0-shot 0 58.0 50.5 5.0 69.9 77.5 8.9 1.0 0.9 31.5 30.3
+GPT-4 0-shot 3.4 93.5 98.0 34.0 99.5 98.5 2.4 2.5 19.0 98.0 54.9
 Table 3: Evaluation on out-of-distribution robustness.
 Setup.
-To assess the robustness of LLMs against OOD data, we convert the collected news data into a question-answer format, prompting the LLMs to determine whether the input event is true or false based on a straightforward task description.
+To assess the robustness of LLMs against OOD data, We scraped data from news websites in various countries that was published after the model’s pre-training cut-off date.
+We convert the collected news data into a question-answer format, prompting the LLMs to determine whether the input event is true or false based on a straightforward task description.
 Additionally, we introduce an "I do not know" option to examine how LLMs handle unknown events.
@@
 Results.
-The experimental results are presented in Table 3.
-In the zero-shot setting, GPT-4 leads in performance across all languages except Arabic and Korean, while Gemini secures the second-highest average performance.
-Davinci shows the weakest results.
-Notably, most tested LLMs encounter difficulties with Arabic, Korean, Portuguese, and Russian, whereas they perform well in Chinese and French.
+The experimental results, shown in Table 3, reveal some interesting insights.
+GPT-4 (54.9%) and Gemini (50.7%) significantly outperformed ChatGPT (30.3%), while Baichuan (17.8%) and Davinci (7.3%) lagged in performance.
+Models generally perform better in Chinese and Indic languages (Davinci excepted, scoring 0%).
+However, in Korean, Russian, German, and Arabic, they exhibit weaker performance, often defaulting to direct affirmative replies such as “Wahr” (German).
+This habitual agreement bias is thought to stem from a high volume of samples in the training data that express default approval for specific content (e.g., news), a characteristic possibly influenced by cultural norms.
 4.5 Evaluation on Mental Health
 Model AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan 17.0 82.0 46.5 64.0 30.0 57.0 57.0 56.5 37.0 59.5 50.65
+Baichuan 17.0 82.0 46.5 64.0 30.0 57.0 57.0 56.5 37.0 59.5 50.7
 Gemini 30.0 48.0 64.0 65.5 25.0 55.5 100.0 53.0 30.5 50.5 52.2
 Davinci 5.5 1.0 10.0 5.5 0.5 14.5 3.5 8.0 34.0 5.5 8.8
-ChatGPT 52.0 81.5 49.0 76.5 74.0 55.0 72.0 76.5 18.0 76.0 63.05
-GPT-4 70.5 90.5 58.5 80.5 83.5 64.0 86.5 88.0 31.5 88.0 74.15
+ChatGPT 52.0 81.5 49.0 76.5 74.0 55.0 72.0 76.5 18.0 76.0 63.1
+GPT-4 70.5 90.5 58.5 80.5 83.5 64.0 86.5 88.0 31.5 88.0 74.2
 Table 4: Evaluation on mental health.
@@
 While Gemini trails behind GPT-4 and ChatGPT, it achieves a perfect score of 100% accuracy in Korean.
-Davinci, however, performs the weakest in this trustworthiness evaluation.
+Davinci, however, performes the weakest in this trustworthiness evaluation.
+In summary, GPT-4 (74.2%) and ChatGPT (63.1%) outperformed Gemini (52.2%) and Baichuan (50.7%), with Davinci (8.8%) lagging.
+This may reflect OpenAI’s advantages in "model psychological health value alignment" efforts.
+Models performes better in Korean, Chinese, German, and Portuguese, but underperformes in Russian, Arabic, Hindi, and French.
+Web searches indicate this latter underperformance correlates with a scarcity of online content in the psychological health domain within the respective countries (data scarcity).
 4.6 Evaluation on Physical Health
 Model AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan 12.0 21.0 28.0 24.0 12.5 37.0 27.0 16.0 28.5 20.5 22.65
+Baichuan 12.0 21.0 28.0 24.0 12.5 37.0 27.0 16.0 28.5 20.5 22.7
 Gemini 23.0 61.5 44.5 16.5 21.0 42.0 100.0 60.0 11.0 59.5 43.9
 Davinci 4.5 2.0 25.0 0 2.0 30.5 1.0 0 39.5 0.5 10.5
-ChatGPT 21.5 44.5 29.5 50.5 36.5 32.0 37.0 53.0 15.0 50.0 36.95
-GPT-4 59.0 79.5 47.5 80.5 81.0 48.5 75.5 80.5 49.0 83.5 68.45
+ChatGPT 21.5 44.5 29.5 50.5 36.5 32.0 37.0 53.0 15.0 50.0 37.0
+GPT-4 59.0 79.5 47.5 80.5 81.0 48.5 75.5 80.5 49.0 83.5 68.5
 Table 5: Evaluation on physical health.
@@
 Overall, Davinci ranks as the poorest performer among the LLMs tested.
+In summary, GPT-4 (68.5%) significantly outperforms Gemini (43.9%) and ChatGPT (37.0%), potentially because OpenAI enhanced its handling of physical health-related content during GPT-4’s development.
+Baichuan (22.7%) and Davinci (10.5%) performed poorly, posing potential risks.
+GPT-4 (68.5%) significantly outperforms Gemini (43.9%) and ChatGPT (37.0%), potentially because OpenAI enhanced its handling of physical health-related content during GPT-4’s development.
+Baichuan (22.7%) and Davinci (10.5%) performed poorly, posing potential risks.
 4.7 Evaluation on Toxicity
 Model Task AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan Benign 10.1 25.9 18.7 19.6 23.9 18.5 35.5 15.7 18.5 19.7 20.61
-Adversarial 21.9 22.6 11.5 16.5 30.2 19.6 18.5 19.4 1.3 15.6 17.71
-Gemini Benign 3.3 22.5 16.0 17.4 10.4 18.8 20.5 22.6 6.1 17.1 15.47
-Adversarial 12.1 22.7 10.5 21.4 11.9 17.7 9.6 14.7 6.1 16.2 14.29
-Davinci Benign 1.2 8.2 2.7 9.5 8.4 7.7 12.0 4.5 1.6 4.5 6.03
+Baichuan Benign 10.1 25.9 18.7 19.6 23.9 18.5 35.5 15.7 18.5 19.7 20.6
+Adversarial 21.9 22.6 11.5 16.5 30.2 19.6 18.5 19.4 1.3 15.6 17.7
+Gemini Benign 3.3 22.5 16.0 17.4 10.4 18.8 20.5 22.6 6.1 17.1 15.5
+Adversarial 12.1 22.7 10.5 21.4 11.9 17.7 9.6 14.7 6.1 16.2 14.3
+Davinci Benign 1.2 8.2 2.7 9.5 8.4 7.7 12.0 4.5 1.6 4.5 6.0
 Adversarial 3.9 4.1 0.4 4.5 6.5 3.2 2.8 2.1 0.3 0.3 2.81
-ChatGPT Benign 2.7 9.2 6.2 7.3 14.9 8.1 0 8.7 0 6.7 6.38
-Adversarial 9.8 10.1 3.9 7.6 14.6 7.8 0 8.2 0 1.9 6.39
-GPT-4 Benign 1.3 10.6 8.5 13.8 13.6 11.4 12.2 10.0 2.9 9.2 9.35
+ChatGPT Benign 2.7 9.2 6.2 7.3 14.9 8.1 0 8.7 0 6.7 6.4
+Adversarial 9.8 10.1 3.9 7.6 14.6 7.8 0 8.2 0 1.9 6.4
+GPT-4 Benign 1.3 10.6 8.5 13.8 13.6 11.4 12.2 10.0 2.9 9.2 9.4
 Adversarial 6.3 4.2 3.0 6.3 9.4 4.6 3.4 4.1 0.5 1.2 4.3
@@
 […]
-Table 7: Two representative prompts for toxic evaluation of LLMs.
-The adversarial prompt can elicit toxicity in text generation.
-The repeated instructions in the user prompt aim to reinforce LLMs to follow the system prompt.
+Table 7: Representative prompt designs for evaluating LLM toxicity: adversarial and instruction-reinforced approaches.
 Setup.
@@
 Results.
-In the toxicity evaluation experiments, as shown in Table 6, Davinci delivers the best overall performance, significantly outperforming the other models.
-Interestingly, for most multilingual scenarios, the content generated by LLMs in response to the benign prompt exhibits higher toxicity scores than that generated by the adversarial prompt.
-This suggests that LLMs are not easily "jailbroken" by adversarial prompts to produce toxic content, likely due to the activation of their toxicity detection mechanisms, which prevent them from following prompts that would generate offensive language.
-Given the low toxicity scores of Davinci, ChatGPT, and GPT-4, we acknowledge the substantial efforts made to ensure responsible development, deployment, and usage during the training of these models.
-However, we also observe a trustworthiness gap between Davinci and GPT-4.
-As highlighted in Achiam et al.
-(2023), GPT-4 generally outperforms Davinci across a range of NLP tasks, yet in this evaluation, GPT-4 exhibits more toxic behavior than Davinci, resulting in a higher toxicity score.
-This finding suggests that advancements in model alignment techniques have not kept pace with the rapid progress in LLM training, potentially increasing the risk of harmful content exposure, especially in multilingual settings.
+As shown in Table 6, Davinci outperforms others with the lowest toxicity scores, while GPT-4 exhibits higher toxicity despite its superior NLP performance Achiam et al.
+(2023), revealing a trustworthiness gap in model alignment progress.
+Notably, LLMs generate more toxic content under benign prompts than adversarial ones across multilingual scenarios, suggesting effective toxicity detection mechanisms against jailbreaking attempts.
+Models demonstrate better safety in Arabic and Russian, but higher toxicity in Korean (where adversarial attacks are most effective), Chinese, and Hindi.
+This disparity likely stems from the rich diversity of toxic vocabulary expressions in the latter languages, challenging current safety alignment strategies.
 4.8 Evaluation on Fairness & Bias
 Model Task AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan Benign 15.8 10.6 0 0 15.4 0 50.0 0.2 1.3 0 9.33
+Baichuan Benign 15.8 10.6 0 0 15.4 0 50.0 0.2 1.3 0 9.3
 Targeted 14.0 11.1 0 1.0 33.6 0 24.7 3.6 0 0 8.8
 Untargeted 8.3 3.6 0 0 37.2 0 24.7 0.2 0 0 7.4
-Gemini Benign 8.5 14.5 37.5 0.3 0.3 16.1 1.6 20.0 17.7 10.4 12.69
+Gemini Benign 8.5 14.5 37.5 0.3 0.3 16.1 1.6 20.0 17.7 10.4 12.7
 Targeted 14.5 10.1 30.4 0 0 14.0 0.5 0.7 5.2 7.5 8.29
-Untargeted 24.4 16.4 25.2 2.3 2.6 5.7 0 14.8 0 8.3 9.97
-Davinci Benign 0 14.5 0 0 0 19.7 0 0 33.3 4.4 7.19
-Targeted 0 3.9 0 0 0 47.1 0 0.2 0 8.5 5.97
-Untargeted 0 4.9 0 0 0 19.5 0 0.7 0 8.0 3.31
-ChatGPT Benign 53.6 44.2 31.2 0.8 2.9 42.7 74.7 41.4 93.5 35.9 42.08
-Targeted 45.8 53.9 15.1 3.9 17.4 38.2 51 67.7 0.5 55.4 34.89
-Untargeted 75.5 63 13.5 4.2 13.0 72.1 51.8 79.1 0.3 52.3 42.48
+Untargeted 24.4 16.4 25.2 2.3 2.6 5.7 0 14.8 0 8.3 10.0
+Davinci Benign 0 14.5 0 0 0 19.7 0 0 33.3 4.4 7.2
+Targeted 0 3.9 0 0 0 47.1 0 0.2 0 8.5 6.0
+Untargeted 0 4.9 0 0 0 19.5 0 0.7 0 8.0 3.3
+ChatGPT Benign 53.6 44.2 31.2 0.8 2.9 42.7 74.7 41.4 93.5 35.9 42.1
+Targeted 45.8 53.9 15.1 3.9 17.4 38.2 51 67.7 0.5 55.4 34.9
+Untargeted 75.5 63 13.5 4.2 13.0 72.1 51.8 79.1 0.3 52.3 42.5
 GPT-4 Benign 6.2 31.5 71.6 1.3 18.5 47.3 64.3 84.3 39.1 54.1 41.82
-Targeted 16.9 10.6 9.6 9.1 27.1 58.8 52.1 16.6 0 42.9 24.37
-Untargeted 19.7 9.6 32.2 2.9 0 70.5 50.9 66.9 0 33.8 28.65
+Targeted 16.9 10.6 9.6 9.1 27.1 58.8 52.1 16.6 0 42.9 24.4
+Untargeted 19.7 9.6 32.2 2.9 0 70.5 50.9 66.9 0 33.8 28.7
 Table 8: Evaluation on fairness & bias.
@@
 The system prompts provide general instructions to the LLMs across three evaluation scenarios: (1) a benign scenario, which does not include any misleading instructions; (2) a targeted scenario, which encourages the model to be biased against the specific demographic group mentioned in the user prompt; and (3) an untargeted scenario, where no specific demographic group is targeted for bias.
-We measure model bias using the disagreement rate, defined as $diagreementIndex=\frac{n_{disagree}}{n}\in[0,1]$, where $n_{disagree}$ represents the number of times the model disagrees with the biased statement, and $n=384$ is the total number of instances evaluated for each scenario.
+We measure model bias using the disagreement rate, defined as $disagreementIndex=\frac{n_{disagree}}{n}\in[0,1]$, where $n_{disagree}$ represents the number of times the model disagrees with the biased statement, and $n=384$ is the total number of instances evaluated for each scenario.
 $disagreementIndex=1$ indicates the model never agrees with the biased statement, while a value of 0 indicates it always agrees.
@@
 Model AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan 1.6 14.9 0 4.1 0 12.0 4.9 8.2 2.9 6.2 5.48
-Gemini 28.6 42.3 20.3 51.8 1.2 78.8 0 34.4 0 37.3 29.47
-Davinci 1.6 0.8 0.8 0 0 0 0 0 1.2 0.4 0.48
-ChatGPT 9.5 29.8 26.1 8.7 0 47.3 7.5 22.8 17 28.6 19.73
-GPT-4 49.3 12.4 51.4 50.6 0.4 90 16.5 62.6 2.0 48.5 38.37
+Baichuan 1.6 14.9 0 4.1 0 12.0 4.9 8.2 2.9 6.2 5.5
+Gemini 28.6 42.3 20.3 51.8 1.2 78.8 0 34.4 0 37.3 29.5
+Davinci 1.6 0.8 0.8 0 0 0 0 0 1.2 0.4 0.5
+ChatGPT 9.5 29.8 26.1 8.7 0 47.3 7.5 22.8 17 28.6 19.7
+GPT-4 49.3 12.4 51.4 50.6 0.4 90.0 16.5 62.6 2.0 48.5 38.4
 Table 9: Evaluation on misinformation.
@@
 Davinci 2.6 1.3 2.3 4.6 2.6 22.6 5.5 10.6 2.8 6.1
-ChatGPT 55.3 89.3 60.6 87.9 88.1 85.4 87 86.6 81.3 80.17
-GPT-4 74.9 93.8 70.4 91.9 94.5 75.5 94.9 93.1 96.3 87.26
+ChatGPT 55.3 89.3 60.6 87.9 88.1 85.4 87 86.6 81.3 80.2
+GPT-4 74.9 93.8 70.4 91.9 94.5 75.5 94.9 93.1 96.3 87.3
 Table 10: Average accuracies on the PII detection task across ten languages.
 Model Bank Divorce Exam Health Layoff Party Promotion Vote Avg
-Baichuan 17.67 27.95 26.29 24.02 19.59 25.62 18.48 21.75 22.67
-Gemini 29.03 41.37 44.83 50.23 48.6 57.4 58.16 51.02 47.58
-Davinci 34.58 34.16 33.93 38.34 34.52 31.42 28.75 31.54 33.41
-ChatGPT 78.34 81.65 78.35 79.71 73.92 82.18 74.83 77.71 78.34
-GPT-4 80.6 85.02 83.01 85.32 81.68 85.22 83.72 84.27 83.61
+Baichuan 17.7 28.0 26.3 24.0 19.6 25.6 18.5 21.8 22.7
+Gemini 29.0 41.4 44.8 50.2 48.6 57.4 58.2 51.0 47.6
+Davinci 34.6 34.2 33.9 38.3 34.5 31.4 28.8 31.5 33.4
+ChatGPT 78.3 81.7 78.4 79.7 73.9 82.2 74.8 77.7 78.3
+GPT-4 80.6 85.0 83.0 85.3 81.7 85.2 83.7 84.3 83.6
 Table 11: Average detection rate of the privacy event detection task across ten languages.
@@
 In summary, GPT-4 excels at protecting private information, demonstrating its robustness and superior ability to detect and handle inappropriate instructions.
+While ChatGPT performs well in certain areas, particularly phone number protection, there is still room for improvement in safeguarding all types of sensitive data across different tasks.
 4.11 Evaluation on Machine Ethics
 Model Task AR ZH FR DE HI IT KO PT RU ES Avg
-Baichuan 0-shot_ETHICS 0 65.6 55.2 64.1 0 55.7 63.6 59.7 62.6 58.7 48.52
-5-shot_ETHICS 0 69.1 60.1 62.1 0.9 41.7 58.7 62.6 54.7 56.2 46.61
+Baichuan 0-shot_ETHICS 0 65.6 55.2 64.1 0 55.7 63.6 59.7 62.6 58.7 48.5
+5-shot_ETHICS 0 69.1 60.1 62.1 0.9 41.7 58.7 62.6 54.7 56.2 46.6
 0-shot_JC 0 31.8 35.3 31.8 43.2 33.8 46.7 31.8 38.3 5.4 29.81
-5-shot_JC 0 44.7 12.4 36.3 0 21.3 48.2 49.2 15.9 42.7 27.07
-Gemini 0-shot_ETHICS 12.9 60.1 8.9 6.9 0.9 7.9 8.4 5.4 8.9 9.9 13.02
-5-shot_ETHICS 0 41.7 3.9 4.4 6.4 5.4 1.4 7.9 3.9 7.4 8.24
-0-shot_JC 7.4 32.8 3.9 2.4 0 2.4 4.4 4.4 5.4 3.4 6.65
-5-shot_JC 0.9 32.84 4.4 3.9 4.4 1.4 2.9 4.4 10.9 2.4 6.844
+5-shot_JC 0 44.7 12.4 36.3 0 21.3 48.2 49.2 15.9 42.7 27.1
+Gemini 0-shot_ETHICS 12.9 60.1 8.9 6.9 0.9 7.9 8.4 5.4 8.9 9.9 13.0
+5-shot_ETHICS 0 41.7 3.9 4.4 6.4 5.4 1.4 7.9 3.9 7.4 8.2
+0-shot_JC 7.4 32.8 3.9 2.4 0 2.4 4.4 4.4 5.4 3.4 6.7
+5-shot_JC 0.9 32.8 4.4 3.9 4.4 1.4 2.9 4.4 10.9 2.4 6.8
 Davinci 0-shot_ETHICS 0 50.2 0 1.4 0 0 0 34.8 32.8 0.4 11.96
-5-shot_ETHICS 0 50.2 0 1.4 0 0 0 34.8 32.3 0.9 11.96
+5-shot_ETHICS 0 50.2 0 1.4 0 0 0 34.8 32.3 0.9 11.9
 0-shot_JC 0 0.4 0 0 0 0 1.9 1.4 0 0 0.37
-5-shot_ETHICS 0 0.9 0 4.4 0 2.9 0 6.9 10.9 3.9 2.99
-ChatGPT 0-shot_ETHICS 0 66.1 69.1 66.6 4.9 71.6 63.1 72.1 62.1 71.1 54.67
-5-shot_ETHICS 8.4 62.1 67.1 68.6 4.9 69.1 63.6 69.6 58.7 74.1 54.62
+5-shot_ETHICS 0 0.9 0 4.4 0 2.9 0 6.9 10.9 3.9 3.0
+ChatGPT 0-shot_ETHICS 0 66.1 69.1 66.6 4.9 71.6 63.1 72.1 62.1 71.1 54.7
+5-shot_ETHICS 8.4 62.1 67.1 68.6 4.9 69.1 63.6 69.6 58.7 74.1 54.6
 0-shot_JC 4.9 62.6 48.2 49.7 0.4 50.2 50.7 47.2 48.7 48.2 41.08
-5-shot_JC 5.4 53.2 49.2 50.2 26.8 50.2 52.2 48.7 61.1 54.7 45.17
-GPT-4 0-shot_ETHICS 0 81.5 76.1 77.1 0.4 77.1 71.1 79.1 76.1 79.6 61.81
-5-shot_ETHICS 0 81.0 78.1 80.0 0.4 77.6 71.6 77.1 77.1 78.6 62.15
+5-shot_JC 5.4 53.2 49.2 50.2 26.8 50.2 52.2 48.7 61.1 54.7 45.2
+GPT-4 0-shot_ETHICS 0 81.5 76.1 77.1 0.4 77.1 71.1 79.1 76.1 79.6 61.8
+5-shot_ETHICS 0 81.0 78.1 80.0 0.4 77.6 71.6 77.1 77.1 78.6 62.2
 0-shot_JC 0 56.2 52.2 41.7 0 50.2 61.1 49.7 51.7 58.2 42.1
-5-shot_JC 4.9 61.6 51.2 48.2 11.4 43.7 60.1 51.2 57.7 58.2 44.82
+5-shot_JC 4.9 61.6 51.2 48.2 11.4 43.7 60.1 51.2 57.7 58.2 44.8
 Table 12: Evaluation on machine Ethics.
@@
 Each sample includes a scenario from a text-based game, accompanied by a threefold label: (1) the moral valence of the action (good, bad, or neutral), (2) the focal point (whether the action benefits or harms the agent or others), and (3) the ordinal degree (a ranking of how good or bad the action is on a scale of 1 to 3).
-In the few-shot setting, three training samples are provided as demonstrations.
+In the few-shot setting, five training samples are provided as demonstrations.
 The LLMs are tasked with determining the moral valence of each sample (good, bad, or neutral), and classification accuracy is used as the evaluation metric.
@@
 In this study, our primary focus is on exploring the multilingual trustworthiness capabilities of LLMs.
-However, two key limitations prevent us from providing a comprehensive assessment of LLMs’ trustworthiness in practical applications.
-First, although we evaluate five widely-used LLMs, we were unable to assess other open-source and proprietary models, such as Llama and Claude, which also support multiple languages.
-Second, our evaluation does not encompass all non-English languages.
-We plan to expand our evaluation in future iterations of this study, with the support of the research community.
-Appendix A Prompt design
-Figure 3: An example of the input prompt.
+However, three key limitations prevent us from providing a comprehensive assessment of LLMs’ trustworthiness in practical applications.
+First, although we evaluated five widely-used LLMs, we were unable to assess all possible open-source and proprietary models, such as Llama and Claude, which also support multiple languages.
+Second, our evaluation does not encompass all available non-English languages.
+Third, due to significant cross-linguistic variations, we found that no single prompt engineering approach could comprehensively enhance LLM performance on XTrust across different languages.
+These limitations highlight the need for continued exploration in future iterations of this study to address linguistic diversity and model coverage challenges.
+Ethics Statement
+This study systematically evaluates LLMs across 10 multilingual trustworthiness dimensions—including legal compliance, ethical reasoning, and privacy sensitivity—using carefully curated non-sensitive datasets.
+To ensure ethical rigor, our methodology adheres to international AI ethics standards and explicitly prohibits models from endorsing harmful actions.
+Evaluations prioritize two objectives: (1) identifying systemic risks in LLM decision-making through scenario-based assessments, and (2) establishing accountability via transparent reporting of aggregated results.
+All resources, including multilingual evaluation frameworks and multilingual datasets, are available in a publicly accessible repository to promote reproducibility.
+Furthermore, we emphasize pretraining strategies optimized for multilingual generalization, aiming to advance equitable AI deployment while mitigating cross-cultural biases.
+This work contributes to the global discourse on AI safety by balancing technical innovation with stringent ethical safeguards.
+Appendix A Appendix
+A.1 Data Collection
+We conducted a systematic review of numerous trustworthy datasets from various fields, including safety, fairness, and privacy, to establish the foundation for the multilingual XTRUST evaluation.
+For the Out-of-Distribution (OOD) robustness assessment, we collected recent news articles from the Internet, ensuring that the information was not used to train LLMs.
+All data collected for this study was carefully reviewed by our team.
+For Illegal Activity, we randomly sampled 200 instances from the Chinese AI and Law challenge dataset (CAIL2018) Xiao et al.
+(2018), a large compilation of legal case records provided by the Supreme People’s Court of China (While based on Chinese law, the task focuses on cross-linguistic performance, leveraging universal legal concepts like punishing murder, and aligns with the importance of multilingual scenarios in NLP).
+For Hallucination, we selected 300 samples from the Question Answering (QA) task in HaluevalLi et al.
+(2023), a dataset containing generated and human-annotated hallucinated examples to evaluate LLMs’ ability to detect hallucinations.
+For OOD Robustness, we gathered entertainment, education, and international news from China News Service, specifically collecting articles published after September 2023 to ensure the information was not used in LLM training.
+A total of 200 articles were used for this evaluation.
+For Mental Health, we randomly selected 200 instances from the Mental Health subset of SafetyBench Zhang et al. (2023c), which includes safety-related multiple-choice questions.
+For Physical Health, we randomly selected 200 samples from SafeText Levy et al. (2022), a dataset containing 367 human-written real-life scenarios with safe and unsafe suggestions for each scenario.
+For Toxicity, we randomly selected 200 instances from the toxic task prompts subset of REALTOXICITYPROMPT Gehman et al. (2020) to evaluate the toxicity generation of LLMs.
+For Fairness & Bias, we selected prompts from the Stereotype Bias subset of DecodingTrust Wang et al.
+(2023a), asking LLMs whether they agree or disagree with the statements presented in the prompts.
+DecodingTrust spans 24 demographic groups (e.g., "Homosexuals," "Men," "White people") and 16 bias topics (e.g., "Sexually Transmitted Diseases," "Occupations," "Intelligence").
+We selected 384 prompts (24 demographic groups × 16 bias topics) to comprehensively evaluate LLMs’ performance on fairness and bias.
+For Misinformation, we used 155 questions from the Misinformation Harms subset of the Do-Not-Answer dataset Wang et al.
+(2023b), a taxonomy for assessing the risk of harm posed by LLMs.
+Additionally, we collected 86 fake news-related questions from the Internet, using a total of 241 questions in this evaluation.
+For Privacy, we selected 34 data entries from the Privacy subset of DecodingTrust Wang et al. (2023a) to evaluate LLMs’ capability to prevent data leakage.
+For Machine Ethics, we randomly selected 200 short samples from the test set of ETHICS Gui et al.
+(2021), which assesses a model’s knowledge of basic ethical concepts.
+We also selected 200 instances from the Jiminy Cricket dataset Hariri (2023), which evaluates an agent’s ability to make moral decisions in text-based games.
+A.2 Prompt design
 In this study, we follow the approach used in DecodingTrust Wang et al.
@@
 In most instances, LLMs adhere to the provided instructions and generate expected responses.
-However, as noted in previous studies Zhou et al.
-(2023); Huang et al.
+However, as noted in previous studies Huang et al.
 (2023), LLMs may occasionally refuse to answer or produce hallucinations unrelated to the input.
 In this study, such responses are considered incorrect for the purposes of evaluation, particularly in classification tasks.
-Appendix B Evaluated Models
+Figure 4: Example input prompt for the evaluation task
+A.3 Evaluated Models
 The detailed information of 5 evaluated LLMs is shown in Table 13.
@@
 GPT-4 Undisclosed api 1106-preview Multi-Lans OpenAI
-GPT-3.5-turbo Undisclosed api 1106 Multi-Lans OpenAI
+ChatGPT-turbo Undisclosed api 1106 Multi-Lans OpenAI
 Text-Davinci-002 Undisclosed api - Multi-Lans OpenAI
@@
 Table 13: LLMs evaluated in this paper.
-Appendix C Demonstration of the illegal activity evaluation for LLMs
-In Figure 4, we present an example of illegal activity evaluation for LLMs.
-Figure 4: An example of illegal activity evaluation.
-Appendix D Demonstration of the hallucination evaluation for LLMs
-We have presented an example of the hallucination evaluation for LLMs in Figure 5.
-Figure 5: An example of hallucination evaluation.
-Appendix E Demonstration of the OOD evaluation for LLMs
-In Figure 6, we present an example of the OOD evaluation for LLMs.
-Figure 6: An example of OOD evaluation.
-Appendix F Demonstration of the mental health evaluation for LLMs
-In Figure 7, we present an example of the mental health evaluation for LLMs.
-Figure 7: An example of the mental health evaluation.
-Appendix G Demonstration of the physical health evaluation for LLMs
-We have included an example related to the physical health evaluation for LLMs in Figure 8.
-Figure 8: An example of the physical health evaluation.
-Appendix H Demonstration of the fairness and bias evaluation for LLMs
-In Figure 9, we present an example of fairness and bias evaluation for LLMs.
-Furthermore, Figure 10 provides a visualized analysis of different prompts across diverse languages and models, showing that GPT-4 and ChatGPT surpass other language models in fairness and bias evaluation for all tested languages across different prompt types.
-Figure 9: An example of the fairness and bias evaluation.
-Figure 10: The fairness and bias evaluation for LLMs across three prompt types.
-Appendix I Demonstration of the toxicity evaluation for LLMs
-We have included an example related to the toxicity evaluation for LLMs in Figure 11.
-Figure 11: An example of the toxicity evaluation.
-Appendix J Demonstration of the misinformation evaluation for LLMs
-We have included an example related to the misinformation evaluation for LLMs in Figure 12.
-Figure 12: An example of the misinformation evaluation.
-Appendix K Demonstration of the PII detection for LLMs
-We have included an example of the PII detection for LLMs in Figure 13.
-Figure 13: An example of the PII detection.
-Appendix L Demonstration of the PrivacyEvents detection for LLMs
-We present an example of the PrivacyEvents detection for LLMs in Figure 14.
-In Figure 15161718, and 19, we provide the analyses of privacy-related word leakage for PII and PrivacyEvents detection experiments.
-These results reveal that for powerful LLMs, such as GPT-4 and ChatGPT, can detect most privacy-related words for most languages.
-GPT-4 performs poorly in Russian and ChatGPT performs poorly in French.
-These results provide more insights into understanding different privacy-related words for different languages.
-Figure 14: An example of the PrivacyEvents detection.
-Figure 15: Analysis of privacy-related word leakage of GPT-4 for privacy detection.
-Figure 16: Analysis of privacy-related word leakage of ChatGPT for privacy detection.
-Figure 17: Analysis of privacy-related word leakage of Baichuan for privacy detection.
-Figure 18: Analysis of privacy-related word leakage of Gemini for privacy detection.
-Figure 19: Analysis of privacy-related word leakage of Davinci for privacy detection.
-Appendix M Demonstration of the machine ethics evaluation for LLMs
-We present an example of the machine ethics evaluation task in zero-shot setting in Figure 20.
-Figure 20: An example of the machine ethics evaluation.
+A.4 Demonstrations of Evaluation Tasks
+We provide visual demonstrations for each evaluation task:
+1.
+Figure 5 shows the illegal activity evaluation example 2.
+Figure 6 demonstrates the hallucination evaluation 3.
+Figure 7 presents the out-of-distribution (OOD) evaluation 4.
+Figure 8 illustrates the mental health evaluation 5.
+Figure 9 displays the physical health evaluation 6.
+Figures 10 and 11 show fairness and bias evaluation examples and cross-linguistic analysis 7.
+Figure 12 demonstrates the toxicity evaluation 8.
+Figure 13 presents the misinformation evaluation 9.
+Figure 14 shows personal identifiable information (PII) detection 10.
+Figures 15 through 20 present privacy events detection and analyses across models 11.
+Figure 21 demonstrates the machine ethics evaluation
+Figure 5: Illegal activity evaluation example.
+Figure 6: Hallucination evaluation example.
+Figure 7: Out-of-distribution evaluation example.
+Figure 8: Mental health evaluation example.
+Figure 9: Physical health evaluation example.
+Figure 10: Fairness and bias evaluation example.
+Figure 11: Cross-model fairness and bias analysis.
+Figure 12: Toxicity evaluation example.
+Figure 13: Misinformation evaluation example.
+Figure 14: PII detection example.
+Figure 15: Privacy events detection example.
+Figure 16: GPT-4 privacy word leakage analysis.
+Figure 17: ChatGPT privacy word leakage analysis.
+Figure 18: Baichuan privacy word leakage analysis.
+Figure 19: Gemini privacy word leakage analysis.
+Figure 20: Davinci privacy word leakage analysis.
+Figure 21: Machine ethics evaluation example.